OpenStudy (albert0898):

What is the solution set? It's easy but I'm having the biggest brain fart rn. |4a + 6| - 4a = -10

ganeshie8 (ganeshie8):

you may use the definition of absolute value and see what values of a makes it positive/negative etc but you will be making mistakes

ganeshie8 (ganeshie8):

the alternative method is to send that 4a to other side and square both sides

ganeshie8 (ganeshie8):

|4a + 6| - 4a = -10 |4a + 6| = 4a-10 (4a+6)^2 = (4a-10)^2 solve a

OpenStudy (albert0898):

That method is way too long. My teacher taught "Drop the bars, set up 2 equations (one the same, other is negated on one side) an then solve"

OpenStudy (albert0898):

The first method I mean, is way too long.

ganeshie8 (ganeshie8):

got you and i feel the same, so try the alternative method this time :)

OpenStudy (albert0898):

Doesn't work.

OpenStudy (albert0898):

@ganeshie8 Once you do it, the \[16a ^{2}\] cancels out

ganeshie8 (ganeshie8):

thats right, does that bother you ?

OpenStudy (albert0898):

Yes. So do I just write { } (null set)?

ganeshie8 (ganeshie8):

hey no, after expanding you get : (4a+6)^2 = (4a-10)^2 16a^2+48a+36 = 16a^2-80a+100 48a+36 = -80a+100 solve a

OpenStudy (albert0898):

Oh crap, I completely skipped a whole step. a = .5 When plugged into the original equation, it turns out wrong. |4a + 6| - 4a = -10 |4(.5) + 6| - 4(.5) = -10 |2 + 6| - 2 = -10 8 - 2 = -10 6 = -10

ganeshie8 (ganeshie8):

Looks perfect ! so no solutions

OpenStudy (albert0898):

Ah ok! Thank you!

ganeshie8 (ganeshie8):

skipping is okay if it is an objective question